# Simple homotopy types of even dimensional manifolds

### v2

@@ 3.1. Realising via $h$-cobordisms
If $(W;M,N)$ is an $h$-cobordism, then we will use the isomorphism $\pi_{1}(i_{M})$ to identify $\pi_{1}(M)$ with $\pi_{1}(W)$, and regard $\tau(W,M)\in\operatorname{Wh}(\pi_{1}(W))$ as an element of $\operatorname{Wh}(\pi_{1}(M))$. Here is the complete statement of the $s$-cobordism theorem, for closed manifolds. It is due to Smale, Barden, Mazur, Stallings, Kirby-Siebenmann, and Freedman-Quinn [Smale-h-cob, Barden, Mazur:1963-1, Stallings-Tata, Kirby-Siebenmann:1977-1, FQ].
@@
Let $M$ be a closed, $\operatorname{CAT}$ $n$-manifold with $\pi_{1}(M)\cong G$, satisfying 1.1.
@@
See [KPR-survey, Theorem 3.5] for details of the proof of part (b) in the case $n=4$. In this case there is an extra subtlety that does not occur for $n\geq 5$, because one has to check that $M^{\prime}$ has the same fundamental group as $M$.
Recall that $\mathcal{I}_{n}(G,w)\coloneqq\{y+(-1)^{n+1}\overline{y}\mid y\in\operatorname{Wh}(G,w)\}\leq\operatorname{Wh}(G,w)$.
@@
Let $M$ be a closed, $\operatorname{CAT}$ $n$-manifold with $\pi_{1}(M)\cong G$ and orientation character $w\colon G\rightarrow\left\{\pm 1\right\}$, satisfying 1.1. For every $x\in\mathcal{I}_{n}(G,w)$ there exists a closed, $\operatorname{CAT}$ $n$-manifold $N$ and a homotopy equivalence $f\colon N\to M$ induced by an $h$-cobordism between $M$ and $N$ such that $\tau(f)=x$.
@@
Let $x\in\mathcal{I}_{n}(G,w)$ and write $x=-y+(-1)^{n}\overline{y}$ for some $y\in\operatorname{Wh}(G,w)$. Apply Theorem 3.1 to obtain an $h$-cobordism $(W;M,N)$ from $M$ to some $n$-manifold $N$ with $\tau(W,M)=y$. If $f$ is the homotopy equivalence induced by $W$, then $\tau(f)=-y+(-1)^{n}\overline{y}=x$ by Proposition 2.38 (b). ∎
@@
Let $M$ and $N$ be closed, $\operatorname{CAT}$ $n$-manifolds satisfying 1.1. Suppose $M$ has $\pi_{1}(M)\cong G$ and orientation character $w\colon G\rightarrow\left\{\pm 1\right\}$. If there is a homotopy equivalence $f\colon N\rightarrow M$ such that $\tau(f)\in\mathcal{I}_{n}(G,w)$, then there exists a closed, $\operatorname{CAT}$ $n$-manifold $P$ that is simple homotopy equivalent to $N$ and $h$-cobordant to $M$.
@@ 3.2. Realising via the surgery exact sequence
We recall the surgery exact sequence. Let $M$ be a closed $\operatorname{CAT}$ $n$-dimensional manifold with $\pi_{1}(M)=G$ and orientation character $w\colon G\to\{\pm 1\}$, for some $n\geq 4$, satisfying 1.1.
@@
The homotopy structure set of $M$, denoted $\mathcal{S}^{h}(M)$, is by definition the set of pairs $(N,f)$, where $N$ is a closed $\operatorname{CAT}$ $n$-manifold and $f\colon N\rightarrow M$ is a homotopy equivalence, considered up to $h$-cobordism over $M$. That is, $[N,f]=[N^{\prime},f^{\prime}]\in\mathcal{S}^{h}(M)$ if and only if there is an $h$-cobordism $(W;N,N^{\prime})$, with inclusion maps $i\colon N\to W$ and $i^{\prime}\colon N^{\prime}\to W$, together with a map $F\colon W\to M$ such that $F\circ i=f$ and $F\circ i^{\prime}=f^{\prime}$.
@@
We can similarly define the simple homotopy structure set $\mathcal{S}^{s}(M)$ to be the set of pairs $(N,f)$, where $N$ is a closed $\operatorname{CAT}$ $n$-manifold and $f\colon N\rightarrow M$ is a simple homotopy equivalence, considered up to $s$-cobordism over $M$.
@@
The Browder-Novikov-Sullivan-Wall surgery exact sequence for $x\in\{h,s\}$ is as follows [Wall-SCM, FQ]; see also [Luck-basic-intro-surgery, Theorem 5.12], [Freedman-book-surgery-chapter].
[…]
@@
Here, $\mathcal{N}(M\times[0,1],M\times\{0,1\})$ and $\mathcal{N}(M)$ denote the sets of the normal bordism classes of degree one normal maps over $M\times[0,1]$ and $M$ respectively. These sets do not depend on the decoration $x$.
@@
The groups $L_{n}^{x}(\mathbb{Z}G,w)$ are the surgery obstruction groups. Elements of $L_{n}^{h}(\mathbb{Z}G,w)$ are represented by nonsingular Hermitian forms over finitely generated free $\mathbb{Z}G$-modules for $n$ even, and by nonsingular formations over finitely generated free $\mathbb{Z}G$-modules for $n$ odd, with the involution on $\mathbb{Z}G$ determined by $w$. See e.g. [Ranicki-80-I]. In the case of $L_{n}^{s}(\mathbb{Z}G,w)$ the forms/formations are also required to be based and simple.
@@
The maps labelled $\sigma_{x}$ are the surgery obstruction maps. For the definition of $\sigma_{s}$, we take a degree one normal map $(f,b)\colon N\to M$, perform surgery below the middle dimension to make the map $[n/2]$-connected, and then produce the based, simple form or formation (for $n$ even or odd respectively) of the surgery kernel in the middle dimension(s), to obtain an element of $L_{n}^{s}(\mathbb{Z}G,w)$. To define the map $\sigma_{h}$ we perform the same procedure, and then forget the data of the bases, to obtain an element of $L_{n}^{h}(\mathbb{Z}G,w)$. One of the main theorems of surgery  [Wall-SCM],[Luck-basic-intro-surgery] is that the maps $\sigma_{x}$, for $x\in\{h,s\}$, are well-defined.
@@
The map $W_{x}$ is the Wall realisation map. Given a $z\in L_{n+1}^{x}(\mathbb{Z}G,w)$ and $[M_{0},f_{0}]\in\mathcal{S}^{x}(M)$, Wall realisation produces a new element $[M_{1},f_{1}]\in\mathcal{S}^{x}(M)$ together with a degree one normal bordism between $(M_{0},f_{0})$ and $(M_{1},f_{1})$ whose surgery obstruction equals $z$. This determines an action of $L_{n+1}^{x}(\mathbb{Z}G,w)$ on $\mathcal{S}^{x}(M)$, and the map $W_{x}$ is defined by acting on the equivalence class of the identity map, $[M,\operatorname{Id}_{M}]\in\mathcal{S}^{x}(M)$.
The $s$ and $h$ decorated $L$ groups are related by the exact sequence:
[…]

### v3

@@ 2.5. The Whitehead torsion of a homotopy equivalence between manifolds
2.6. Milnor’s definition of Whitehead torsion
So far we defined the Whitehead torsion of a (chain) homotopy equivalence. Now we recall Milnor’s definition [Mi66, Sections 3–4], which applies in broader settings. We will use this more general notion (only) in the surgery theoretic arguments of Section 3.2. We slightly alter the presentation compared to [Mi66], so that we do not need to manipulate (s-)bases directly, and we will see that these alterations do not change the result.
We fix a group $G$, and assume that all modules are finitely generated left $\mathbb{Z}G$-modules.
Definition 2.39.
A module $X$ is called stably free if $X\oplus(\mathbb{Z}G)^{k}$ is free for some $k\geq 0$, and an s-basis (or stable basis) of $X$ is the data of an integer $k$ and a basis of $X\oplus(\mathbb{Z}G)^{k}$ for some $k$. An s-based module will mean a stably free module equipped with an s-basis.
If $X$ is free, then a basis of $X$ is also an s-basis. A trivial module has a unique basis, which determines the canonical s-basis for it.
Definition 2.40.
Suppose that $X$ and $Y$ are s-based modules, with $X\oplus(\mathbb{Z}G)^{k}$ and $Y\oplus(\mathbb{Z}G)^{l}$ free and based, and $f\colon X\rightarrow Y$ is an isomorphism. If $r\geq\max(k,l)$, then the s-bases determine bases for $X\oplus(\mathbb{Z}G)^{r}$ and $Y\oplus(\mathbb{Z}G)^{r}$ (by adding the standard bases of $(\mathbb{Z}G)^{r-k}$ and $(\mathbb{Z}G)^{r-l}$), and $f\oplus\operatorname{Id}_{(\mathbb{Z}G)^{r}}$ is an isomorphism between them. The matrix of this isomorphism in the given bases represents an element of $\operatorname{GL}(\mathbb{Z}G)$, and hence $\operatorname{Wh}(G)$. This element is independent of the choice of $r$, and it is called the Whitehead torsion of the isomorphism $f$, denoted $\tau(f)$.
Remark 2.41.
Two s-bases of a stably free module $X$ are said to be equivalent if the isomorphism $\operatorname{Id}_{X}$ between them has vanishing Whitehead torsion. As the Whitehead torsion is additive under composition, the Whitehead torsion of an isomorphism $f\colon X\rightarrow Y$ only depends on the equivalence classes of the s-bases of $X$ and $Y$. Hence in all statements about s-based modules, it is actually sufficient to equip the modules with an equivalence class of s-bases.
Definition 2.42.
Suppose that
$0\rightarrow X\rightarrow Y\rightarrow Z\rightarrow 0$
is a short exact sequence of s-based $\mathbb{Z}G$-modules. Since a stably free module (in particular, $Z$) is projective, such a sequence splits, and a splitting determines an isomorphism $Y\cong X\oplus Z$ of s-based modules. Since changing the splitting changes the isomorphism by an element of $E(\mathbb{Z}G)$, the Whitehead torsion of this isomorphism is independent of the choice of splitting (cf. [Mi66, Section 2]), and is defined to be the Whitehead torsion of the short exact sequence. (In [Mi66] this is denoted by $[xz/y]$, where $x,y,z$ are the (s-)bases of the modules.)
Suppose that $C_{*}$ is a chain complex of s-based modules, and assume that the modules $H_{i}:=H_{i}(C_{*})$ are also s-based. In such a situation we will always assume that all but finitely many of the modules $C_{i}$ and $H_{i}$ are trivial and equipped with the canonical s-basis. By [Mi66, Section 4] the submodules $B_{i}\leq Z_{i}\leq C_{i}$ of boundaries and cycles are also stably free (and all but finitely many of them are trivial).
Definition 2.43.
We define the Whitehead torsion of an s-based chain complex $C_{*}$ with s-based homology as follows. Choose s-bases for the modules $B_{i}$ and $Z_{i}$ such that all but finitely many of them are trivial and equipped with the canonical s-bases, and the s-bases of the remaining ones are chosen arbitrarily. Denote the Whitehead torsions of the short exact sequences
$0\rightarrow Z_{i}\rightarrow C_{i}\rightarrow B_{i-1}\rightarrow 0\quad\quad\quad\text{and}\quad\quad\quad 0\rightarrow B_{i}\rightarrow Z_{i}\rightarrow H_{i}\rightarrow 0$
by $\gamma_{i}$ and $\zeta_{i}$ in $\operatorname{Wh}(G)$ respectively. Then only finitely many of the $\gamma_{i}$ and $\zeta_{i}$ are non-zero, and we define
$\tau(C_{*}):=\sum_{i}(-1)^{i}(\gamma_{i}+\zeta_{i})\in\operatorname{Wh}(G).$
One checks that changing the s-basis of $B_{i}$ or $Z_{i}$ has no effect on the sum, so the Whitehead torsion of $C_{*}$ does not depend on the choice of s-bases of $B_{i}$ and $Z_{i}$. By choosing an s-basis for $Z_{i}$ such that $\zeta_{i}$ vanishes (denoted by $b_{i}h_{i}$ in [Mi66]) we see that this definition coincides with the one given in [Mi66, Sections 3–4].
Definition 2.44.
The Whitehead torsion of an acyclic s-based chain complex is defined by taking the canonical s-basis for all of the (trivial) homology groups and using Definition 2.43.
If $f\colon C_{*}\to D_{*}$ is a chain homotopy equivalence between finitely generated, free, based chain complexes, then its mapping cone $\mathscr{C}(f)$ is an acyclic, free, based chain complex. By [Co73, §16], $\tau(f)$ of Definition 2.11 and $\tau(\mathscr{C}(f))$ of Definition 2.44 coincide in $\operatorname{Wh}(G)$, showing that Milnor’s definition of torsion generalises the one we gave earlier. In fact, if we replace the mapping cone $\mathscr{C}(f)$ in Definition 2.11 with any acyclic, free, based chain complex, the resulting notion of Whitehead torsion would coincide with that given by Definition 2.44 [Co73, §15–16].
The next three lemmas follow directly from the definitions.
Lemma 2.45.
Suppose that $f\colon X\rightarrow Y$ is an isomorphism of s-based modules, and $C_{*}$ is an acyclic chain complex such that $C_{i}=X$, $C_{i-1}=Y$ ((with the given s-bases)), $d_{i}=f$, and for $j\neq i,i-1$, $C_{j}$ is trivial with the canonical s-basis. Then $\tau(C_{*})=(-1)^{i}\tau(f)$.
Lemma 2.46.
Suppose that $0\rightarrow X\rightarrow Y\rightarrow Z\rightarrow 0$ is a short exact sequence of s-based modules, and is regarded as an acyclic chain complex $C_{*}$, with $C_{i+1}=X$, $C_{i}=Y$, and $C_{i-1}=Z$. Then its Whitehead torsion is $(-1)^{i}\tau(C_{*})$.
Lemma 2.47.
Suppose that $C_{*}$ is an s-based chain complex with trivial differentials, so that $H_{i}=C_{i}$ for all $i$. Assume that $H_{i}$ is equipped with the same s-basis as $C_{i}$ for each $i$. Then $\tau(C_{*})=0$.
The behaviour of torsion in short exact sequences is very important.
Theorem 2.48 ([Mi66, Theorem 3.2]).
Suppose that $^{\prime}_{*}, CC_{*}$ and $C^{\prime\prime}_{*}$ are s-based chain complexes with s-based homology. Suppose also that $0\rightarrow C^{\prime}_{*}\rightarrow C_{*}\rightarrow C^{\prime\prime}_{*}\rightarrow 0$ is a short exact sequence of chain complexes such that the short exact sequence $0\rightarrow C^{\prime}_{i}\rightarrow C_{i}\rightarrow C^{\prime\prime}_{i}\rightarrow 0$ has vanishing Whitehead torsion for all $i$. Let $D_{*}$ be the associated homological long exact sequence, regarded as an acyclic s-based chain complex, with $D_{3i+2}=H^{\prime}_{i}$, $D_{3i+1}=H_{i}$ and $D_{3i}=H^{\prime\prime}_{i}$. Then $\tau(C_{*})=\tau(C^{\prime}_{*})+\tau(C^{\prime\prime}_{*})+\tau(D_{*})$.
The next lemma shows how we can use the Whitehead torsion to specify an s-basis of a homology group.
Lemma 2.49.
Suppose that $C_{*}$ is an s-based chain complex and $i$ is an integer such that $H_{j}$ is s-based for $j\neq i$ and $H_{i}$ is stably free. Then for every $x\in\operatorname{Wh}(G)$ there is an s-basis of $H_{i}$ ((unique up to equivalence)) such that $\tau(C_{*})=x$.
Proof.
Choose an arbitrary s-basis of $H_{i}$ and let $y\in\operatorname{Wh}(G)$ be the Whitehead torsion of $C_{*}$ with respect to that choice. There is an $r\geq 0$ such that the chosen s-basis of $H_{i}$ determines a basis of $H_{i}\oplus(\mathbb{Z}G)^{r}$ and the element $(-1)^{i}(x-y)\in\operatorname{Wh}(G)$ can be represented by an $f\in\operatorname{GL}_{r}(\mathbb{Z}G)$. Then a suitable s-basis of $H_{i}$ can be obtained by applying $\operatorname{Id}_{H_{i}}\oplus f$ to the original basis of $H_{i}\oplus(\mathbb{Z}G)^{r}$. ∎
Note that, even when $H_{i}$ is free, this construction only determines an s-basis of $H_{i}$ (as opposed to a basis), because it cannot be guaranteed that a given element of the Whitehead group is represented by an automorphism of a free module of a fixed rank. The following special case is used in the definition of the surgery obstruction (see [Wall-SCM, Lemma 2.3]).
Definition 2.50.
Suppose that $C_{*}$ is an s-based chain complex and $i$ is an integer such that $H_{j}$ is trivial for $j\neq i$ and $H_{i}$ is stably free. The distinguished s-basis of $H_{i}$ is defined (up to equivalence) by equipping $H_{j}$ with the canonical s-basis for $j\neq i$ and applying Lemma 2.49 with $x=0$.
If $H_{i}$ is trivial too, then it also has the canonical s-basis. However, in general the distinguished s-basis differs from the canonical one.
Lemma 2.51.
Suppose that $C_{*}$ is an acyclic s-based chain complex and $\tau(C_{*})$ is its Whitehead torsion as in Definition 2.44. Then the transition matrix between the canonical and the distinguished s-basis of $H_{i}$ represents $(-1)^{i+1}\tau(C_{*})\in\operatorname{Wh}(G)$.
Finally, we can use Milnor’s definition of Whitehead torsion in the geometric setting.
Definition 2.52.
Suppose that $Y$ is a finite CW complex with $\pi_{1}(Y)=G$, $X$ is a subcomplex of $Y$ (possibly empty), and $H_{i}(Y,X;\mathbb{Z}G)$ is an s-based $\mathbb{Z}G$-module for all $i$. Then the Whitehead torsion of a pair $(Y,X)$, denoted $\tau(Y,X)$, is defined to be $\tau(C_{*}(Y,X;\mathbb{Z}G))$ (with the basis of $C_{i}(Y,X;\mathbb{Z}G)$ given by the $i$-cells of $Y$ not in $X$).
Lemma 2.53.
Suppose that $\pi_{1}(Y)=G$, $f\colon X\rightarrow Y$ is a cellular homotopy equivalence between finite CW complexes, and $H_{i}(\operatorname{Cyl}_{f},X;\mathbb{Z}G)=0$ is equipped with the canonical s-basis for all $i$. Then $\tau(\operatorname{Cyl}_{f},X)=\tau(f)$.
Proof.
Since $C_{*}(\operatorname{Cyl}_{f},X)\cong\mathscr{C}(f_{*})$, we have $\tau(\operatorname{Cyl}_{f},X)=\tau(C_{*}(\operatorname{Cyl}_{f},X))=\tau(\mathscr{C}(f_{*}))=\tau(f_{*})=\tau(f)$ by the definitions and the observation below Definition 2.44. ∎
We can now check that $\tau(Y,X)$ is invariant under simple homotopy equivalence, in a suitable sense.
Lemma 2.54.
Suppose that $(Y,X)$ and $(Y^{\prime},X^{\prime})$ are pairs with identifications $\pi_{1}(Y)=\pi_{1}(Y^{\prime})=G$, and $H_{i}(Y,X;\mathbb{Z}G)$ and $H_{i}(Y^{\prime},X^{\prime};\mathbb{Z}G)$ are s-based. Suppose also that there are simple homotopy equivalences $f\colon X\rightarrow X^{\prime}$ and $F\colon Y\rightarrow Y^{\prime}$ commuting with the inclusions $X\rightarrow Y$ and $X^{\prime}\rightarrow Y^{\prime}$, such that $\pi_{1}(F)=\operatorname{Id}_{G}$ and the induced isomorphisms $H_{i}(Y,X;\mathbb{Z}G)\rightarrow H_{i}(Y^{\prime},X^{\prime};\mathbb{Z}G)$ respect the s-bases. Then $\tau(Y,X)=\tau(Y^{\prime},X^{\prime})$.
Proof.
Consider the short exact sequences
$0\rightarrow C_{*}(\operatorname{Cyl}_{f},X)$ $\rightarrow C_{*}(\operatorname{Cyl}_{F},Y)\rightarrow C_{*}(\operatorname{Cyl}_{F},Y\cup\operatorname{Cyl}_{f})\rightarrow 0$
$0\rightarrow C_{*}(\operatorname{Cyl}_{f},X^{\prime})$ $\rightarrow C_{*}(\operatorname{Cyl}_{F},Y^{\prime})\rightarrow C_{*}(\operatorname{Cyl}_{F},Y^{\prime}\cup\operatorname{Cyl}_{f})\rightarrow 0$
of chain complexes with $\mathbb{Z}G$ coefficients. All of these chain complexes are acyclic, and we choose the canonical s-bases on their homologies. Since $f$, $F$, and the inclusions $X^{\prime}\rightarrow\operatorname{Cyl}_{f}$ and $Y^{\prime}\rightarrow\operatorname{Cyl}_{F}$ are simple (see Proposition 2.30), by Lemma 2.53 and Theorem 2.2 we have that
$\tau(\operatorname{Cyl}_{f},X)=\tau(\operatorname{Cyl}_{F},Y)=\tau(\operatorname{Cyl}_{f},X^{\prime})=\tau(\operatorname{Cyl}_{F},Y^{\prime})=0,$
and hence by Theorem 2.48 it follows that
$\tau(\operatorname{Cyl}_{F},Y\cup\operatorname{Cyl}_{f})=\tau(\operatorname{Cyl}_{F},Y^{\prime}\cup\operatorname{Cyl}_{f})=0.$
Next, consider the short exact sequences
$0\rightarrow C_{*}(Y,X)$ $\rightarrow C_{*}(\operatorname{Cyl}_{F},\operatorname{Cyl}_{f})\rightarrow C_{*}(\operatorname{Cyl}_{F},Y\cup\operatorname{Cyl}_{f})\rightarrow 0$
$0\rightarrow C_{*}(Y^{\prime},X^{\prime})$ $\rightarrow C_{*}(\operatorname{Cyl}_{F},\operatorname{Cyl}_{f})\rightarrow C_{*}(\operatorname{Cyl}_{F},Y^{\prime}\cup\operatorname{Cyl}_{f})\rightarrow 0.$
Recall that $H_{*}(\operatorname{Cyl}_{F},Y\cup\operatorname{Cyl}_{f})=H_{*}(\operatorname{Cyl}_{F},Y^{\prime}\cup\operatorname{Cyl}_{f})=0$ have both been equipped with the canonical s-basis. We have $H_{*}(\operatorname{Cyl}_{F},\operatorname{Cyl}_{f})\cong H_{*}(Y^{\prime},X^{\prime})$, so we can equip the former with the s-basis given on the latter. Then the homological long exact sequences associated to the above short exact sequences have vanishing Whitehead torsion (for the first sequence, we use that the induced isomorphisms $H_{i}(Y,X;\mathbb{Z}G)\rightarrow H_{i}(Y^{\prime},X^{\prime};\mathbb{Z}G)$ respect the s-bases). So by Theorem 2.48, applied twice, we see that $\tau(Y,X)=\tau(\operatorname{Cyl}_{F},\operatorname{Cyl}_{f})=\tau(Y^{\prime},X^{\prime})$. ∎
@@ 3.1. Realising via $h$-cobordisms
If $(W;M,N)$ is an $h$-cobordism, then we use the isomorphism $\pi_{1}(i_{M})$ to identify $\pi_{1}(M)$ with $\pi_{1}(W)$, and regard $\tau(W,M)\in\operatorname{Wh}(\pi_{1}(W))$ as an element of $\operatorname{Wh}(\pi_{1}(M))$. Here is the complete statement of the $s$-cobordism theorem, for closed manifolds. It is due to Smale, Barden, Mazur, Stallings, Kirby-Siebenmann, and Freedman-Quinn [Smale-h-cob, Barden, Mazur:1963-1, Stallings-Tata, Kirby-Siebenmann:1977-1, FQ].
@@
Let $M$ be a closed, $\operatorname{CAT}$ $n$-manifold with $\pi_{1}(M)\cong G$, with $\operatorname{CAT}$ as in 1.1.
@@
For $n\geq 5$, see [RS82, p. 90] for a discussion of part (b). In the case $n=4$, [KPR-survey, Theorem 3.5] gives details of the proof of part (b). In this case there is an extra subtlety that does not occur for $n\geq 5$, because one has to check that $M^{\prime}$ has the same fundamental group as $M$.
Recall that $\mathcal{I}_{n}(G,w):=\{y+(-1)^{n+1}\overline{y}\mid y\in\operatorname{Wh}(G,w)\}\leq\operatorname{Wh}(G,w)$.
@@
Let $M$ be a closed, $\operatorname{CAT}$ $n$-manifold with $\pi_{1}(M)\cong G$ and orientation character $w\colon G\rightarrow\left\{\pm 1\right\}$, with $\operatorname{CAT}$ as in 1.1. For every $x\in\mathcal{I}_{n}(G,w)$ there exists a closed, $\operatorname{CAT}$ $n$-manifold NN and a homotopy equivalence $f\colon N\to M$ induced by an $h$-cobordism between $M$ and $N$ such that $\tau(f)=x$.
@@
Let $x\in\mathcal{I}_{n}(G,w)$ and write $x=-y+(-1)^{n}\overline{y}$ for some $y\in\operatorname{Wh}(G,w)$. Apply Theorem 3.1 to obtain an $h$-cobordism $(W;M,N)$ from $M$ to some $n$-manifold NN with $\tau(W,M)=y$. If $f$ is the homotopy equivalence induced by $W$, then $\tau(f)=-y+(-1)^{n}\overline{y}=x$ by Proposition 2.38 (b). ∎
@@
Let $M$ and $N$ be closed, $\operatorname{CAT}$ $n$-manifolds, with $\operatorname{CAT}$ as in 1.1. Suppose $M$ has $\pi_{1}(M)\cong G$ and orientation character $w\colon G\rightarrow\left\{\pm 1\right\}$. If there is a homotopy equivalence $f\colon N\rightarrow M$ such that $\tau(f)\in\mathcal{I}_{n}(G,w)$, then there exists a closed, $\operatorname{CAT}$ $n$-manifold $P$ that is simple homotopy equivalent to $N$ and $h$-cobordant to $M$.
@@ 3.2. Realising via the surgery exact sequence
We recall the surgery exact sequence. Let $M$ be a closed $\operatorname{CAT}$ $n$-dimensional manifold with $\pi_{1}(M)=G$ and orientation character $w\colon G\to\{\pm 1\}$, for some $n\geq 4$, with $\operatorname{CAT}$ as in 1.1.
@@
The homotopy structure set of $M$, denoted $\mathcal{S}^{\mathrm{h}}(M)$, is by definition the set of pairs $(N,f)$, where $N$ is a closed $\operatorname{CAT}$ $n$-manifold and $f\colon N\rightarrow M$ is a homotopy equivalence, considered up to $h$-cobordism over $M$. That is, $N,f]=[N^{\prime},f^{\prime}]\in\mathcal{S}^{\mathrm{h}}(M)$ if and only if there is an $h$-cobordism $(W;N,N^{\prime})$, with inclusion maps $i\colon N\to W$ and $i^{\prime}\colon N^{\prime}\to W$, together with a map $F\colon W\to M$ such that $F\circ i=f$ and $F\circ i^{\prime}=f^{\prime}$.
@@
We can similarly define the simple homotopy structure set $\mathcal{S}^{\mathrm{s}}(M)$ to be the set of pairs $(N,f)$, where $N$ is a closed $\operatorname{CAT}$ $n$-manifold and $f\colon N\rightarrow M$ is a simple homotopy equivalence, considered up to $s$-cobordism over $M$. Such a pair $(N,f)$ also represents an element of $\mathcal{S}^{\mathrm{h}}(M)$, inducing a forgetful map $\mathcal{S}^{\mathrm{s}}(M)\rightarrow\mathcal{S}^{\mathrm{h}}(M)$.
@@
The main tool for studying the structure sets $\mathcal{S}^{x}(M)$, for $x\in\{{\mathrm{h}},{\mathrm{s}}\}$, is the Browder-Novikov-Sullivan-Wall surgery exact sequence [Wall-SCM, FQ]; see also [Luck-basic-intro-surgery, Theorem 5.12], [Freedman-book-surgery-chapter]. The two versions of the sequence, and the commutative diagram formed by taking the forgetful maps between them, are as follows.
${\mathcal{N}(M\times[0,1],M\times\{0,1\})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\sigma_{\mathrm{s}}}$${L_{n+1}^{\mathrm{s}}(\mathbb{Z}G,w)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{W_{\mathrm{s}}}$$\scriptstyle{F}$${\mathcal{S}^{\mathrm{s}}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\eta_{\mathrm{s}}}$${\mathcal{N}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\sigma_{\mathrm{s}}}$${L_{n}^{\mathrm{s}}(\mathbb{Z}G,w)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{F}$${\mathcal{N}(M\times[0,1],M\times\{0,1\})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\sigma_{\mathrm{h}}}$${L_{n+1}^{\mathrm{h}}(\mathbb{Z}G,w)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{W_{\mathrm{h}}}$${\mathcal{S}^{\mathrm{h}}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\eta_{\mathrm{h}}}$${\mathcal{N}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\sigma_{\mathrm{h}}}$${L_{n}^{\mathrm{h}}(\mathbb{Z}G,w)}$ (3.1)
Next we explain the remaining terms and maps in these sequences.
  • $\bullet$
@@
    The sets of the normal bordism classes of degree one normal maps over $M\times[0,1]$ and $M$ are denoted $\mathcal{N}(M\times[0,1],M\times\{0,1\})$ and $\mathcal{N}(M)$ respectively. In the former case, the normal map is required to restrict to a $\operatorname{CAT}$ isomorphism over $M\times\{0,1\}$. These sets do not depend on the decoration $x$.
  • $\bullet$
@@
    The groups $L_{n}^{x}(\mathbb{Z}G,w)$ are the surgery obstruction groups. Elements of $L_{n}^{\mathrm{h}}(\mathbb{Z}G,w)$ are represented by nonsingular Hermitian forms over finitely generated free $\mathbb{Z}G$-modules for $n$ even, and by nonsingular formations over finitely generated free $\mathbb{Z}G$-modules for $n$ odd, with the involution on $\mathbb{Z}G$ determined by $w$. See e.g. [Ranicki-80-I] for the definitions. In the case of $L_{n}^{\mathrm{s}}(\mathbb{Z}G,w)$ the forms/formations are also required to be based and simple. The map $F$ is given by forgetting the bases.
  • $\bullet$
@@
    The maps labelled $\sigma_{x}$ are the surgery obstruction maps. For the definition of $\sigma_{\mathrm{s}}$, we take a degree one normal map $(f,b)\colon N\to M$, perform surgery below the middle dimension to make the map $[n/2]$-connected, and then produce the based, simple form or formation (for $n$ even or odd respectively) of the surgery kernel in the middle dimension(s), to obtain an element of $L_{n}^{\mathrm{s}}(\mathbb{Z}G,w)$. To define the map $\sigma_{\mathrm{h}}$ we perform the same procedure, and then forget the data of the bases, to obtain an element of $L_{n}^{\mathrm{h}}(\mathbb{Z}G,w)$. One of the main theorems of surgery [Wall-SCM], [Luck-basic-intro-surgery] is that the maps $\sigma_{x}$ are well-defined.
  • $\bullet$
    We write $\eta_{x}$ for the normal invariant map. It sends $N,f]\in\mathcal{S}^{x}(M)$ to the normal bordism class of $f\colon N\rightarrow M$ covered by a suitable bundle map.
  • $\bullet$
@@
    The map $W_{x}$ is the Wall realisation map. Given $z\in L_{n+1}^{x}(\mathbb{Z}G,w)$ and $M_{0},f_{0}]\in\mathcal{S}^{x}(M)$, Wall realisation produces a new element $M_{0},f_{0}]^{z}=[M_{1},f_{1}]\in\mathcal{S}^{x}(M)$ together with a degree one normal bordism between $(M_{0},f_{0})$ and $(M_{1},f_{1})$ whose surgery obstruction equals $z$. This determines an action of $L_{n+1}^{x}(\mathbb{Z}G,w)$ on $\mathcal{S}^{x}(M)$, and the map $W_{x}$ is defined by acting on the equivalence class of the identity map, $M,\operatorname{Id}_{M}]\in\mathcal{S}^{x}(M)$.
Each term has a natural basepoint, and the sequences are exact sequences of pointed sets. Furthermore, the fibres of $\eta_{x}$ coincide with the orbits of the action of $L_{n+1}^{x}(\mathbb{Z}G,w)$ on $\mathcal{S}^{x}(M)$.
The forgetful maps $F$ fit into the following exact sequence of abelian groups:
$\cdots\to L_{n+1}^{\mathrm{s}}(\mathbb{Z}G,w)\xrightarrow{F}L_{n+1}^{\mathrm{h}}(\mathbb{Z}G,w)\xrightarrow{\psi}\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))\xrightarrow{\varrho}L_{n}^{\mathrm{s}}(\mathbb{Z}G,w)\xrightarrow{F}L_{n}^{\mathrm{h}}(\mathbb{Z}G,w)\to\cdots$ (3.2)